OpenStudy (anonymous):

(3 square root of 5 + square root of 3) (square root of 5 - 7 square root of 3)

OpenStudy (anonymous):

can you put in fourmula form?

OpenStudy (anonymous):

\[(3\sqrt{5}+\sqrt{3})(\sqrt{5}-7\sqrt{3}\]

OpenStudy (anonymous):

should be another parenthese after the squae root of 3

OpenStudy (anonymous):

-83.4 is what i got

OpenStudy (anonymous):

they want u to foil it out

OpenStudy (anonymous):

they dont want a decimal